Please show work for how to find the answer.You are the producer for the television show Actuarial Idol. Each year, 1000 actuarial clubs audition forthe show The probabi
Answer #1

Hello Sir/ Mam

Answer is Option E : 49,000

Given that:

Mean = 20, Variance = 20, hence, standard deviation = 4.472

Now, P(Selection) = 0.2

Total number of team =1000

Hence, the expected number of selected teams = 1000*0.2 = 200

Hence,

Budget = 10(200*20) + (4.472*200*10) = 48944.27 approximately equal to 49,000

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.